Jack Johnson has a high-deductible insurance policy. His deductible is $1,000 and reimbursement is on a 90-10 basis. To date, his charges have totaled $1,689. What amount is due from the insurance company? What amount is due from the patient?
Answer #1

Jack Johnson will have to first pay $1,000 (which the is the amount of deductible) towards the charges before he becomes entitled to insurance amount. The balance amount of $689 will be adjusted in the ratio of 90:10. That is, the insurance company will pay 90% of the balance charges and 10% will be paid by the patient.

Amount Due from the Insurance Company = Balance Amount*90% = 689*90% = $620.10

Amount Due from the Patient = Balance Amount*10% = 689*10% = $68.90